# Real-Time Position-Aware View Synthesis from Single-View Input

### v1

@@ 5.2.3. Results
Quantitative Results: The quantitative results of our light field reconstruction experiments are presented in Table 3. Our proposed PLFNet demonstrates competitive performance compared to state-of-the-art methods, including NoisyLFRecon (Zhou et al., 2021), DGLF (Zhou et al., 2020), DALF (Cun et al., 2019), IR-V (Han and Xiang, 2022), and LFSphereNet (Gond et al., 2023). All models were trained from scratch using identical datasets for a fair comparison. Publicly available implementations were used for the methods by Zhou et al. (2021, 2020) and Han and Xiang (2022), while DALF (Cun et al., 2019) was re-implemented in PyTorch according to the original paper.
@@
Although DGLF and LFSphereNet generate the full light field in a single pass, both DALF and our PLFNet reconstruct each sub-aperture image iteratively. For the Stanford dataset, which offers fewer training samples compared to Lytro Flowers, LFSphereNet delivered the best results across most metrics, with PLFNet showing competitive performance, particularly excelling in SSIM and MS-SSIM. In terms of computational efficiency, PLFNet demonstrated significant improvements as shown in Table 4. Our PLFNet achieved an inference time of 0.0980 seconds for whole LF reconstruction which included forward passes. However, per view inference on average is around 0.006 seconds highlighting its potential for real-time applications.
@@
Qualitative Results: The qualitative evaluation of light field reconstruction is illustrated in Figure 5. We compare the visual performance of our proposed PLFNet against state-of-the-art methods for light field views, focusing on the top right corner sub-aperture view. The first two rows depict reconstruction results for the JPEG-Pleno (Rerabek and Ebrahimi, 2016) dataset. In these challenging scenarios, PLFNet demonstrates superior reconstruction quality with fewer visible artifacts, as evident in the error maps. Compared to LFSphereNet (Gond et al., 2023), which also shows strong performance, PLFNet exhibits more accurate details and reduced error magnitudes, highlighting its robustness and generalization capabilities. The subsequent rows display results for the Stanford Light Field Archive (Raj et al., 2016) dataset. Here, our method consistently outperforms NoisyLFRecon (Zhou et al., 2021), which utilizes nine input images for reconstruction. Despite this advantage, NoisyLFRecon produces noticeable artifacts and blurring in regions with fine structures. In contrast, PLFNet, which relies on only a single input image, achieves sharper reconstructions with more faithful reproduction of scene geometry and textures. This indicates that PLFNet can effectively capture and synthesize complex light field information even with minimal input data.
@@
Overall, the qualitative results corroborate the quantitative findings, showcasing ability of PLFNet to achieve high-quality reconstructions with low error rates. This performance shows potential for practical applications where high-fidelity light field reconstruction from sparse data is required.
@@ 6. Ablation Study
In this section, we present a two-part ablation study to assess the effectiveness of our proposed method, focusing on positional understanding and the impact of different positional embedding techniques. The first part evaluates the performance of PLFNet+ under varying sparse input configurations, demonstrating how the number and location of input images influence view synthesis. The second part investigates the contribution of different positional encoding strategies to the overall reconstruction quality, highlighting the role of each component in our final model.
@@ 6.1. Sparse Input Mode
The first part of the ablation study explores how the performance of PLFNet+ is affected by varying the number and position of input images for view synthesis. We evaluate three configurations: (i) a single central image as input, (ii) a single image from a random location within the cubical volume, and (iii) two images from random locations. These setups allow us to understand how well the model generalizes to sparse input scenarios, particularly in cases where inputs are not centrally aligned.
@@
Table 5 presents the quantitative results of this study. As expected, the single central view input yields the best performance across all metrics, with a PSNR of 28.8023 and SSIM of 0.9009, indicating that the model performs best when provided with a centrally located reference image. The random input configurations, whether with one or two images, exhibit slightly reduced performance, suggesting that while the model can generalize to inputs from random locations, the quality of the synthesized view degrades with less structured input data. Nevertheless, even with these random inputs, PLFNet+ achieves acceptable performance, demonstrating its robustness in handling diverse input scenarios.
| Number of Input Images | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| 2 Images - Random Location | 23.0571 | 0.7444 | 0.8462 | 0.4297 | 0.2782 | 0.2510 |
| 1 Image - Random Location | 22.8429 | 0.7243 | 0.8386 | 0.3524 | 0.3373 | 0.3703 |
| 1 Image - Central View | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 |
6.2. Positional Embedding
In the second part of the ablation study, we analyze the impact of different positional embedding strategies on view synthesis. We compare four configurations: (i) MLP only, (ii) normalization only, (iii) normalization with positional encoding, (iv) normalization with MLP, and (v) the full setup with normalization, positional encoding, and MLP. This analysis aims to determine the contribution of each component in enhancing the model’s understanding of positional information.
The results are summarized in Table 6. The combination of normalization, positional encoding, and MLP consistently outperforms the other configurations, achieving a PSNR of 28.8023 and SSIM of 0.9003. This demonstrates the effectiveness of our final proposed method in capturing positional variations and synthesizing accurate novel views. By contrast, using only normalization or MLP leads to a noticeable drop in performance specially in high frequency regions, underscoring the importance of combining these techniques to achieve optimal results.
The setup with normalization and MLP alone performs better than using MLP in isolation but still falls short of the complete configuration. These findings highlight the critical role of positional encoding and MLP in providing a richer representation of the 3D positional information, which is essential for accurately reconstructing complex novel views.
| Positional Embedding | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| MLP | 23.2578 | 0.7633 | 0.8686 | 0.4063 | 0.2566 | 0.2726 |
| Normalization | 25.2855 | 0.8376 | 0.9256 | 0.5309 | 0.2132 | 0.1680 |
| Norm+Positional Encoding | 24.1945 | 0.7941 | 0.8962 | 0.4502 | 0.2276 | 0.1946 |
| Norm+MLP | 26.4475 | 0.8508 | 0.9421 | 0.6247 | 0.2119 | 0.1647 |
| Norm+PosEnc+MLP | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 |
7. Conclusion
In this paper we presented a novel, lightweight, position-aware network for real-time view synthesis and light field reconstruction, which leverages a position-aware embedding combined with a rendering network to achieve both efficiency and high visual quality. Through the development of PLFNet and its extension, PLFNet+, our approach demonstrates superior synthesis quality, particularly in handling complex translational movements without relying on computationally intensive operations like warping.
Extensive evaluations reveal that PLFNet+ achieves a strong balance between visual fidelity and inference speed, positioning it as a feasible solution for applications requiring real-time performance, such as telepresence and augmented reality. Additionally, our network architecture shows robustness in various scenarios, from narrow to wide baseline settings, making it adaptable to dynamic and interactive environments.
The promising results from ablation studies underscore the effectiveness of our positional embedding design, particularly the integration of normalization, positional encoding, and MLP components. Overall, our work provides a significant step towards more accessible, high-performance view synthesis frameworks, paving the way for broader application in consumer and industrial devices.
@@ Acknowledgements.
The work was supported by the European Joint Doctoral Programme on Plenoptic Imaging (PLENOPTIMA) through the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 956770. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. We thank the High Performance Computing Center North (HPC2N) at Umeå University for providing computational resources and valuable support during test and performance runs.References
@@
- Bae et al. (2021) Kyuho Bae, Andre Ivan, Hajime Nagahara, and In Kyu Park. 2021. 5d light field synthesis from a monocular video. In 2020 25th International Conference on Pattern Recognition (ICPR). IEEE, 7157–7164.
@@
- Chandramouli et al. (2020) Paramanand Chandramouli, Kanchana Vaishnavi Gandikota, Andreas Goerlitz, Andreas Kolb, and Michael Moeller. 2020. A generative model for generic light field reconstruction. IEEE Transactions on Pattern Analysis and Machine Intelligence 44, 4 (2020), 1712–1724.
- Chen et al. (2020) Bin Chen, Lingyan Ruan, and Miu-Ling Lam. 2020. LFGAN: 4D Light Field Synthesis from a Single RGB Image. ACM Trans. Multimedia Comput. Commun. Appl. 16 (2 2020). Issue 1. https://doi.org/10.1145/3366371
@@
- Ivan et al. (2019) Andre Ivan, In Kyu Park, et al. 2019. Synthesizing a 4D spatio-angular consistent light field from a single image. arXiv preprint arXiv:1903.12364 (2019).

### v2

@@ 4.2.3. Results
Quantitative Image Quality: The quantitative results of our light field reconstruction experiments are presented in Table 3. Our proposed PVSNet demonstrates competitive performance compared to state-of-the-art methods, including NoisyLFRecon (Zhou et al., 2021), DGLF (Zhou et al., 2020), DALF (Cun et al., 2019), IR-V (Han and Xiang, 2022), and LFSphereNet (Gond et al., 2023). All models were trained from scratch using identical datasets for a fair comparison. Publicly available implementations were used for the methods by Zhou et al. (2021, 2020) and Han and Xiang (2022), while DALF (Cun et al., 2019) was re-implemented in PyTorch according to the original paper.
@@
Although DGLF and LFSphereNet generate the full light field in a single pass, both DALF and our PVSNet reconstruct each sub-aperture image iteratively. For the Stanford dataset, which offers fewer training samples compared to Lytro Flowers, LFSphereNet delivered the best results across most metrics, with PVSNet showing competitive performance, particularly excelling in SSIM and MS-SSIM. However, we note that while LFSphereNet and the other methods in this comparison were specifically designed for light field reconstruction, PVSNet is optimized for view synthesis within the narrow baseline of light fields. Consequently, PVSNet does not require the blending of sub-aperture views to generate the final synthesized view, setting it apart from the other methods in this comparison.
Quantitative Processing Time: In terms of computational efficiency, PVSNet demonstrated significant improvements as shown in Table 4. Our PVSNet achieved an inference time of 0.0850 seconds for the whole LF reconstruction which included forward passes. However, PVSNet allows for direct querying of any target view within the light field baseline, enabling significantly faster per-view inference, averaging around 0.0079 seconds (126 fps). This efficiency highlights its suitability for real-time applications. In contrast, methods like LFSphereNet and DALF, despite achieving faster overall light field reconstruction times, necessitate the blending of sub-aperture views using per-view disk based blending algorithm (Overbeck et al., 2018). This process requires copying the light field data to the GPU buffer and utilizing shaders to blend the views for final synthesis. This additional step introduces an overhead of 1.301 seconds on an RTX 2070 Super, resulting in considerably slower time of 1.3018 seconds for LFSphereNet and 1.3587 seconds for DALF as shown in the last column of Table 4. Therefore, PVSNet presents a more efficient solution for narrow baseline view synthesis.
| MLP | Normalization | Norm+PosEnc | Norm+MLP | Norm+PosEnc+MLP |
| Positional Embedding | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| MLP | 23.6780 | 0.8267 | 0.8698 | 0.4577 | 0.1521 | 0.1063 |
| Normalization | 25.5882 | 0.8466 | 0.9172 | 0.5673 | 0.1152 | 0.0766 |
| Norm+Positional Encoding | 24.2862 | 0.8105 | 0.8719 | 0.4675 | 0.1376 | 0.0919 |
| Norm+MLP | 26.7322 | 0.8581 | 0.9366 | 0.6422 | 0.1070 | 0.0888 |
| Norm+PosEnc+MLP | 31.2258 | 0.9386 | 0.9793 | 0.8702 | 0.0613 | 0.0275 |
@@
Qualitative Image Quality: The qualitative evaluation of light field reconstruction is illustrated in Figure 7. We compare the visual performance of our proposed PVSNet against state-of-the-art methods for light field views, focusing on the top right corner sub-aperture view. The first two rows depict reconstruction results for the JPEG-Pleno (Rerabek and Ebrahimi, 2016) dataset. In these challenging scenarios, PVSNet demonstrates superior reconstruction quality with fewer visible artifacts, as evident in the error maps. Compared to LFSphereNet (Gond et al., 2023), which also shows strong performance, PVSNet exhibits more accurate details and reduced error magnitudes, highlighting its robustness and generalization capabilities. The subsequent rows display results for the Stanford Light Field Archive (Raj et al., 2016) dataset. Here, our method consistently outperforms NoisyLFRecon (Zhou et al., 2021), which utilizes nine input images for reconstruction. Despite this advantage in quantitative results where the average scores are higher, NoisyLFRecon produces noticeable artifacts and blurring in regions with fine structures. In contrast, PVSNet, which relies on only a single input image, achieves sharper reconstructions with more faithful reproduction of scene geometry and textures. This indicates that PVSNet can effectively capture and synthesize complex light field information even with minimal input data. More visual results can be found in our supplementary materials video, where we show the ability of PVSNet to generate any views within the LF baseline by directly querying the target position.
@@
Overall, the qualitative results corroborate the quantitative findings, showcasing ability of PVSNet to achieve high-quality reconstructions with low error rates. This performance shows potential for practical applications where high-fidelity light field reconstruction from sparse data is required.
@@ 5. Ablation Study
In this section, we present a two-part ablation study to assess the effectiveness of our proposed method, focusing on positional understanding and the impact of different positional embedding techniques. The first part investigates the contribution of different positional embedding strategies to the overall reconstruction quality, highlighting the role of each component in our final model. The second part investigates the effect of modifying the structure of the Rendering Network, focusing on the contribution of Encoder I. Finally, the third part evaluates the performance of PVSNet under varying sparse input configurations, demonstrating how the number and location of input images influence view synthesis.
5.1. Positional Embedding
In the first part of the ablation study, we analyze the impact of different positional embedding strategies on view synthesis. We compare four configurations as shown in Figure 8: (i) MLP only, (ii) normalization only, (iii) normalization with positional encoding, (iv) normalization with MLP, and (v) the full setup with normalization, positional encoding, and MLP. This analysis aims to determine the contribution of each component in enhancing the model’s understanding of positional information. For a fair comparison, we train each variation of the network for same number of epochs.
The results are summarized in Table 5. (i) MLP only: The model using only the MLP for the positional embedding block without any form of normalization or positional encoding exhibits the weakest performance. The lack of effective positional information results in inability of the rendering network to capture high-frequency variations in the scene, leading to a reconstruction that predominantly represents low-frequency features. This is reflected in a significantly lower average PSNR of 23.6780 dB, indicating poor view synthesis quality shown in Figure 9. (ii) Normalization only: When only coordinate normalization is employed, the network performs reasonably well for small baselines, where the scene geometry and appearance differences between the source and target viewpoints are minimal. However, as the baseline increases, the target viewpoint deviates more significantly from the source, leading to greater variations in scene appearance. Consequently, the model struggles to generalize to these larger baselines, yielding a PSNR of 25.5882 dB. This result highlights the limited capacity of normalization alone to effectively handle complex geometric transformations. (iii) Normalization with Positional Encoding: Introducing sinusoidal positional encoding on the normalized coordinates improves ability of the model to account for large geometric variations between source and target views. However, this approach also increases the search space of the rendering network as it must learn to map the given positional embedding vector to a suitable output space. This increased complexity can lead to suboptimal learning, as evidenced by the barbershop scene in Figure 9, where the rendered image exhibits a noticeable color mismatch. This configuration results in a PSNR of 24.2862 dB.
(iv) Normalization with MLP: By combining normalization with an MLP, the model benefits from the ability of the MLP to project the normalized coordinates into a higher-dimensional space. This allows the rendering network to better capture both low and high-frequency variations in the scene, making it more effective at handling large baselines. The model achieves a PSNR of 26.7322 dB, demonstrating a marked improvement over the previous configurations. This result shows the importance of using an MLP to model complex transformations between source and target viewpoints, especially in scenarios involving significant changes in scene appearance.
(v) Normalization, Positional Encoding, and MLP: The final configuration, which incorporates normalization, positional encoding, and an MLP, provides the most comprehensive solution. By applying positional encoding after normalization, the network benefits from a more structured and compact search space, allowing the MLP to more effectively map the positional inputs to a corresponding high dimensional space. This leads to a richer feature representation of the 3D positional information, facilitating the accurate reconstruction of complex novel views. As a result, this configuration achieves the highest PSNR of 31.2258 dB, demonstrating its superior performance in synthesizing novel views.
Overall, the results from this ablation study highlight the significant contributions of each component in enhancing the understanding of positional information. The combination of normalization, positional encoding, and MLP provides the most robust and effective solution for accurate and high-quality view synthesis. This combination allows the model to better represent complex geometric and appearance changes, which is essential for synthesizing realistic novel views from a single input image.
| Encoder I | Encoder II | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS | FPS |
| ResNet | Ours | 31.2258 | 0.9386 | 0.9793 | 0.8702 | 0.0613 | 0.0275 | 62 |
| DINOv2 | Ours | 32.5534 | 0.9529 | 0.9846 | 0.8945 | 0.0412 | 0.0258 | 23 |
| - | Ours | 22.9666 | 0.8126 | 0.9288 | 0.6965 | 0.1251 | 0.0960 | 69 |
| Number of Input Images | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| 2 Images - Random Location | 25.7561 | 0.8723 | 0.9255 | 0.6328 | 0.1021 | 0.0977 |
| 1 Image - Random Location | 25.0533 | 0.8298 | 0.9041 | 0.5379 | 0.1227 | 0.1076 |
| 1 Image - Central View | 31.2258 | 0.9386 | 0.9793 | 0.8702 | 0.0613 | 0.0275 |
5.2. Rendering Network Components
In the second part of the ablation study, we analyze the architectural choices made in the rendering network, particularly the use of a dual-encoder setup. The original architecture consists of two encoder branches: Encoder I: a ResNet-based feature extractor that processes the input image alone, and Encoder II: a custom designed encoder that jointly processes the input image and the positional embedding to generate position aware features. These features are later fused and passed through a decoder to produce the synthesized view.
To understand the contribution of each encoder component, we evaluate three configurations (see Table 6): (i) the baseline setup using ResNet for Encoder I and our custom block for Encoder II, (ii) a variant where ResNet is replaced with DINOv2 (Oquab et al., 2023) to extract more semantically rich features, and (iii) a simplified setup where Encoder I is entirely removed, relying solely on Encoder II.
The results demonstrate that while removing Encoder I simplifies the architecture and achieves the highest FPS (69), it significantly degrades synthesis quality across all metrics, including a PSNR drop of over 8 dB. This indicates that the positional features alone are insufficient for capturing fine appearance details without dedicated image encoding. This supports our architectural choice to include an additional image only encoder.
Replacing ResNet with DINOv2 in Encoder I further improves quality, achieving the highest PSNR (32.55 dB), SSIM (0.95), and lowest perceptual error (LPIPS of 0.0258). However, this comes at a notable computational cost, reducing inference speed from 62 to 23 FPS. Given our real-time rendering objective, we select ResNet as the backbone for Encoder I to maintain a good balance between visual quality and speed.
These findings validate the design choice of using a dual-encoder architecture, where Encoder I provides a strong appearance prior, and Encoder II injects pose specific spatial cues. The two branches complement each other, leading to sharper, more consistent novel views compared to relying solely on position-aware embeddings.
@@ 5.3. Sparse Input Mode
The third part of the ablation study explores how the performance of PVSNet is affected by varying the number and position of input images for view synthesis. We evaluate three configurations: (i) a single central image as input, (ii) a single image from a random location within the cubical volume, and (iii) two images from random locations. These setups allow us to understand how well the model generalizes to sparse input scenarios, particularly in cases where inputs are not centrally aligned.
@@
Table 7 presents the quantitative results of this study. As expected, the single central view input yields the best performance across all metrics, with a PSNR of 31.2258 dB and SSIM of 0.9386, indicating that the model performs best when provided with a centrally located reference image. The random input configurations, whether with one or two images, exhibit reduced performance, suggesting that while the model can generalize to inputs from random locations, the quality of the synthesized view degrades significantly with less structured input data. Nevertheless, even with these random inputs, PVSNet achieves acceptable performance, demonstrating its robustness in handling diverse input scenarios.
6. Conclusion and Discussions
This paper presents PVSNet and its efficient position-aware view synthesis framework, which offers a significant leap in the real-time generation of novel views from a single input image and a target position. Our method demonstrates superior efficiency, achieving real-time performance with a high inference speed, making it particularly suitable for interactive applications like immersive telepresence, 360∘ video with 6DoF and augmented reality (AR). Through extensive testing, PVSNet outperforms several state-of-the-art methods, such as AdaMPI (Han et al., 2022).
Contribution to Existing Solutions: PVSNet enhances the field of view synthesis by introducing an effective solution for real-time inference. Compared to existing methods such as Quark (Flynn et al., 2024) or AdaMPI (Han et al., 2022), which require multi-view inputs or are computationally expensive, PVSNet excels in generating high-quality novel views from a single image. Its efficiency, particularly with large baselines, marks a significant improvement over previous approaches which require multiple images, or struggle with high computational demands or with generating continuous perspectives. While approaches like NViST (Jang and Agapito, 2024) and the use of Multi-Plane Images (MPI) have set the foundation for accurate scene representation, they often face limitations when applied in real-time applications due to their memory and computational overheads. PVSNet addresses these challenges by optimizing both computational efficiency and visual quality, thus supplementing existing solutions and opening up possibilities for their deployment in real-time industrial applications.
Implications and Impact: The outcomes of this research hold substantial implications for both the academic community and industrial applications. For researchers, PVSNet provides a more accessible approach to view synthesis, offering a platform for further innovation in the area of real-time view synthesis. The use of positional embeddings for view synthesis could inspire further exploration of more efficient representations using smaller vision transformer models for real-time scene understanding.
For engineers and industrial applications, the real-time nature of PVSNet makes it a powerful tool for sectors such as remote operations, training simulations, and VR experiences. By providing continuous perspective views with minimal latency, this framework could drastically improve user interactions in these fields. Furthermore, the ability to synthesize novel views from a single input image opens up new possibilities for systems with limited computational resources, where traditional multi-view methods might not be feasible.
However, while the method achieves impressive results, there are still areas for improvement. One limitation is the handling of occlusions and complex geometries in highly dynamic environments when the rendering volume is large for novel view synthesis.
Future Directions: As PVSNet serves as a foundation for real-time position-aware view synthesis, future work can extend this approach by exploring the way to use the latent space generated by the model for other scene understanding tasks, such as depth estimation, segmentation and object detection, to further enhance its applicability in diverse environments. Future research could explore ways to further enhance depth estimation or integrate multi-view information in a hybrid manner, balancing computational efficiency with the ability to handle more complex scene interactions. Additionally, expanding its generalization capabilities to handle a wider range of real-world scenes and augmenting its robustness for large-scale deployment will be crucial steps toward achieving broader industry adoption.
@@ Acknowledgements.
The work was supported by the European Joint Doctoral Programme on Plenoptic Imaging (PLENOPTIMA) through the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 956770, by the Knowledge Foundation, Sweden, with grant number 2019-0251, and by Mid Sweden University internal funding. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. We thank the High Performance Computing Center North (HPC2N) at Umeå University for providing computational resources and valuable support during test and performance runs.References
@@
- Ahn and Kim (2013) Ilkoo Ahn and Changick Kim. 2013. A novel depth-based virtual view synthesis method for free viewpoint video. IEEE Transactions on Broadcasting 59, 4 (2013), 614–626.
@@
- Bertel et al. (2019) Tobias Bertel, Neill DF Campbell, and Christian Richardt. 2019. Megaparallax: Casual 360 panoramas with motion parallax. IEEE transactions on visualization and computer graphics 25, 5 (2019), 1828–1835.
@@
- Chandan et al. (2021) Kishan Chandan, Jack Albertson, Xiaohan Zhang, Xiaoyang Zhang, Yao Liu, and Shiqi Zhang. 2021. Learning to guide human attention on mobile telepresence robots with 360 vision. In 2021 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS). IEEE, 5297–5304.
- Chen and Koltun (2017) Qifeng Chen and Vladlen Koltun. 2017. Photographic image synthesis with cascaded refinement networks. In Proceedings of the IEEE international conference on computer vision. 1511–1520.
- Chen and Williams (1993) Shenchang Eric Chen and Lance Williams. 1993. View interpolation for image synthesis. In Proceedings of the 20th Annual Conference on Computer Graphics and Interactive Techniques (Anaheim, CA) (SIGGRAPH ’93). Association for Computing Machinery, New York, NY, USA, 279–288. https://doi.org/10.1145/166117.166153
@@
- Dima et al. (2020) Elijs Dima, Kjell Brunnström, Mårten Sjöström, Mattias Andersson, Joakim Edlund, Mathias Johanson, and Tahir Qureshi. 2020. Joint effects of depth-aiding augmentations and viewing positions on the quality of experience in augmented telepresence. Quality and User Experience 5 (2020), 1–17.
@@
- Fehn (2004) Christoph Fehn. 2004. Depth-image-based rendering (DIBR), compression, and transmission for a new approach on 3D-TV. In Stereoscopic displays and virtual reality systems XI, Vol. 5291. SPIE, 93–104.
@@
- Flynn et al. (2024) John Flynn, Michael Broxton, Lukas Murmann, Lucy Chai, Matthew DuVall, Clément Godard, Kathryn Heal, Srinivas Kaza, Stephen Lombardi, Xuan Luo, Supreeth Achar, Kira Prabhu, Tiancheng Sun, Lynn Tsai, and Ryan Overbeck. 2024. Quark: Real-time, High-resolution, and General Neural View Synthesis. ACM Trans. Graph. 43, 6, Article 194 (Nov. 2024), 20 pages. https://doi.org/10.1145/3687953
@@
- Jampani et al. (2021) Varun Jampani, Huiwen Chang, Kyle Sargent, Abhishek Kar, Richard Tucker, Michael Krainin, Dominik Kaeser, William T Freeman, David Salesin, Brian Curless, et al. 2021. Slide: Single image 3d photography with soft layering and depth-aware inpainting. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 12518–12527.
@@
- Konrad et al. (2017) Robert Konrad, Donald G Dansereau, Aniq Masood, and Gordon Wetzstein. 2017. Spinvr: towards live-streaming 3d virtual reality video. ACM Transactions on Graphics (TOG) 36, 6 (2017), 1–12.
@@
- Oquab et al. (2023) Maxime Oquab, Timothée Darcet, Théo Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, et al. 2023. Dinov2: Learning robust visual features without supervision. arXiv preprint arXiv:2304.07193 (2023).
@@
- Ranftl et al. (2022) René Ranftl, Katrin Lasinger, David Hafner, Konrad Schindler, and Vladlen Koltun. 2022. Towards Robust Monocular Depth Estimation: Mixing Datasets for Zero-Shot Cross-Dataset Transfer. IEEE Transactions on Pattern Analysis and Machine Intelligence 44, 3 (2022), 1623–1637. https://doi.org/10.1109/TPAMI.2020.3019967
@@
- Sakashita et al. (2023) Mose Sakashita, Hyunju Kim, Brandon Woodard, Ruidong Zhang, and François Guimbretière. 2023. VRoxy: Wide-Area Collaboration From an Office Using a VR-Driven Robotic Proxy. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (San Francisco, CA, USA) (UIST ’23). Association for Computing Machinery, New York, NY, USA, Article 59, 13 pages. https://doi.org/10.1145/3586183.3606743
@@
- Shih et al. (2020) Meng-Li Shih, Shih-Yang Su, Johannes Kopf, and Jia-Bin Huang. 2020. 3d photography using context-aware layered depth inpainting. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 8028–8038.
- Solovev et al. (2023) Pavel Solovev, Taras Khakhulin, and Denis Korzhenkov. 2023. Self-improving multiplane-to-layer images for novel view synthesis. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision. 4309–4318.
@@
- Suhail et al. (2022) Mohammed Suhail, Carlos Esteves, Leonid Sigal, and Ameesh Makadia. 2022. Generalizable patch-based neural rendering. In European Conference on Computer Vision. Springer, 156–174.
@@
- Waidhofer et al. (2022) John Waidhofer, Richa Gadgil, Anthony Dickson, Stefanie Zollmann, and Jonathan Ventura. 2022. PanoSynthVR: Toward Light-weight 360-Degree View Synthesis from a Single Panoramic Input. In 2022 IEEE International Symposium on Mixed and Augmented Reality (ISMAR). IEEE, 584–592.
@@
- Yao et al. (2018) Yao Yao, Zixin Luo, Shiwei Li, Tian Fang, and Long Quan. 2018. MVSNet: Depth Inference for Unstructured Multi-view Stereo. European Conference on Computer Vision (ECCV) (2018).
@@
- Zhang et al. (2024) Jason Y Zhang, Amy Lin, Moneish Kumar, Tzu-Hsuan Yang, Deva Ramanan, and Shubham Tulsiani. 2024. Cameras as rays: Pose estimation via ray diffusion. arXiv preprint arXiv:2402.14817 (2024).
@@
- Zhu et al. (2019) Shiping Zhu, Hao Xu, and Lina Yan. 2019. An improved depth image based virtual view synthesis method for interactive 3D video. IEEE Access 7 (2019), 115171–115180.